OpenStudy (anonymous):

Find the asymptotes of the following function and situation regarding elllas curve: \[f (x)= \frac{ x ^{2} }{ x-3 }\]

OpenStudy (anonymous):

well we know x cannot be 3 because this would make the denominator 0

OpenStudy (anonymous):

there is no horizontal asymptote bcz the degree f denominator is high than numerator . for vertical asymptote set denominator equal to zero x-3=0 x=3

OpenStudy (anonymous):

also a "slant" asymptote if you need it

OpenStudy (anonymous):

just as a side note: if the degree of numerator and denominator is equal then the Horizontal asymptote is just the ratio of the coefficients of the highest degrees .for Example y=(3x+6)/(5x-7) since the degree of x in both numerator and denominator is 1 so horizontal asymptote. should be ratio of the coefficients. y=3/5 is the horizontal asymptote. if you have y=(5x^2+9)/(6x^2+8) since degree of both numerator and denominator is same so y=5/6 is horizontal asymptote . for vertical asymptote you need to solve denominator by setting it equal to zero. if you have y=(x+6)/(x+5) so set denominator equal to zero x+5=0 or x=-5 is vertical asymptote.